# spacing=nonfrench

### v1

@@ 4.3.2 Vulnerable And Perfectly Exploitable
ChaCha20-Poly1305 [30] directly uses the sequence number in its key derivation, which makes it vulnerable to our prefix truncation attack. All messages following the truncated prefix are decrypted to their original plaintext, because the integrity check of the AEAD cipher is only over the ciphertext and the sequence number that has been manipulated by the attacker to match. Under the assumption that the attacker can correctly guess the packet length, the prefix truncation attack always succeeds.
Note that the fault is not with ChaCha20-Poly1305 as an AEAD encryption scheme, but with its integration into the SSH secure channel construction.
4.3.3 Vulnerable, But Not Exploitable
@@
With CTR-EtM, the MAC is computed over the unencrypted length, the sequence number, and the ciphertext. So, removing some packets from the beginning of the channel does not cause a MAC failure, and cryptographically, the attack succeeds. However, CTR uses a block counter initialized to $\mathrm{IV}_{\mathrm{KDF}}$, which is incremented after each block. After prefix truncation, the key stream is desynchronized, so all following ciphertexts are decrypted as pseudo-random packets. Each corrupted packet has a significant probability of causing a critical failure, which eventually stops our attack.
@@
For CTR-EtM, prefix truncation of $k$ blocks (which exactly contain one or more messages) provides a very limited decryption oracle on the ciphertext $c_{1},\ldots,c_{k}$ where $c_{i}:=\mathrm{Enc(\mathrm{IV}_{\mathrm{KDF}}+i)}\oplus p_{i},1\leq i\leq k$. After deleting the first $k$ blocks, MAC verification for the following message of length $l$ blocks will succeed because the length, sequence number, and ciphertext are correct. The blocks $c_{k+1},\ldots,c_{k+l}$ will be decrypted as $p^{\prime}_{j}:=\mathrm{Enc(\mathrm{IV}_{\mathrm{KDF}}+j)}\oplus c_{k+j},1\leq j\leq l$, and processed as a pseudo-random SSH message SC1’. Due to format oracle side channels in SSH at the BPP layer, e.g. the padding length, but also at the protocol layer, e.g. if a message is ignored or triggers a response, the attacker can get some information about the bits in $p^{\prime}_{j}$. This reveals information about the first $l$ key stream blocks, and thus also about $p_{1},\ldots,p_{l}$, potentially leaking confidential information like passwords in user authentication. If processing SC1’ does not cause a critical failure, the attack can even continue further with the decryption of the next message, revealing more about the following key stream and, thus, plaintext. Exploiting this requires a careful study of format oracles in SSH, which is outside the scope of this work.
4.3.4 Vulnerable And Probabilistically Exploitable
@@
Recall that the decryption of the first block is $p_{1}:=\mathrm{Dec}(c_{1})\oplus\mathrm{IV}_{\mathrm{KDF}}$ and for block $i$ it is $p_{i}:=\mathrm{Dec}(c_{i})\oplus c_{i-1}$. We assume the attacker uses prefix truncation to remove blocks $c_{1},\ldots,c_{k}$. The following block $c_{k+1}$ will now be decrypted as $p^{\prime}_{1}:=\mathrm{Dec}(c_{k+1})\oplus\mathrm{IV}_{\mathrm{KDF}}$. We are interested in how SSH implementations process the resulting pseudo-random block $p^{\prime}_{1}$ as the first block in the decrypted packet. Intuitively, it should result in a corrupted packet that causes a critical failure.  [^1] Similarly to CTR-EtM, any format oracle side channel for $p^{\prime}_{1}$ reveals a relationship between $\mathrm{IV}_{\mathrm{KDF}}$ and $p_{k+1}$ via $\mathrm{IV}_{\mathrm{KDF}}\oplus p_{k+1}=c_{k}\oplus p^{\prime}_{1}$, which is a marginal information leak for the (secret) IV given information on $p_{k+1}$, and vice versa. Again, we do not explore this further here.
Surprisingly, there is a significant probability that the attack can continue, although it is highly implementation-dependent. For a corrupted packet, there are three possible outcomes:
@@
    Critically Corrupt: If the corruption is detected at the BPP or application level, e.g., some length field exceeds the packet length, the connection should be closed.
@@
    Evasively Corrupt: If the packet is well-formed (i.e., has valid padding length) but has an unrecognized message ID, an Unimplemented response must be sent, and the connection continues normally [29, Sec. 11.4].
Clearly, the first outcome stops any attack from going forward. However, the second and third outcome may be beneficial for the attacker. We will now present two instructive scenarios for these two outcomes and estimate the success probability of an attack relying on that outcome. Later, we will see some experimental verification for these estimates.
@@
In this scenario, the attacker wants to remove the first message, and the second, corrupted message needs to be functionally preserved but has some format flexibility. For example, the second message might be ServiceAccept (see Section 5.2), which is critical to start user authentication. The encrypted part of the packet looks like this, where $p$ is the padding length, $m$ the message id, and $n$ the service name length:
{bytefield}
@@
Although ServiceAccept may be a lucky case, there are structural reasons for this result: First, SSH messages are often short and can be smaller than a single block. Second, the padding is random and cannot be verified. Third, some messages have redundant fields that implementations ignore (e.g., the service name above).
We experimentally verified that OpenSSH, Dropbear, PuTTY, and libssh allow empty ServiceAccept messages from the server, enabling this attack. At the same time, AsyncSSH is very strict and requires the correct service name.
@@
In this scenario, we assume the attacker wants to remove the first $N>1$ messages and perfectly preserve all of the following messages. Then, the attacker can use prefix truncation to delete the first $N-1$ messages, and take a bet on the $N$-th message to be evasively corrupt.
@@
Because the message is well-formed, it is not rejected at the BPP layer. Furthermore, because the message is unrecognized, the peer must respond with Unimplemented and otherwise ignore it [29, Sec. 11.4], so our attack succeeds.
@@
The probability that a packet is unrecognized depends on the implementation. The attack requires at least one unknown message ID. Through source code review, we identified 43 IDs that are in active use, so we estimate up to 213 unknown message IDs out of $2^{8}$.
In total, assuming a block size of at least 128-bit (i.e., $\ell\geq 16$), we estimate that the success probability of this attack is between $11\cdot 2^{-16}\approx 0.0002$ and $252\cdot 213\cdot 2^{-16}\approx 0.8190$ for vulnerable implementations. Our experiments show success probabilities from 0.0003–0.8383, in good agreement with our analysis (Section 5.2). Increasing the block size increases the lower bound, while the upper bound stays the same.
@@ 5 Breaking SSH Extension Negotiation
While the fact that BPP does not implement a secure channel is troublesome enough, exploiting this vulnerability requires an analysis of the SSH protocol at the application layer.
As our attack achieves prefix truncation, it is only natural to ask which SSH messages can occur at the beginning of a secure channel. Historically, the first messages exchanged are ServiceRequest and ServiceAccept. Removing either causes the connection to go stale, as the client will not begin the user authentication protocol. Then our attack, while cryptographically successful, fails at the application layer.
However, the SSH Extension Negotiation mechanism [9] introduces a new message, ExtInfo, which can occur immediately after NewKeys as the first message on the secure channel. Some of the extensions that can be negotiated are security-relevant, providing a relevant attack surface for our prefix truncation attack and raising its impact.
@@ 5.1 SSH Extension Negotiation
Support for extension negotiation is signaled as part of the KexInit message. To not pose compatibility issues, the structure of the message is not being altered, and the reserved field is not used. Instead, each peer may include an indicator name within the list of key exchange algorithms. The indicator name differs depending on the role of the peer (ext-info-c vs. ext-info-s) to avoid accidental negotiation.
Whenever a peer signals support for extension negotiation, the other side may send an ExtInfo message as the first message after NewKeys. Additionally, the server can send a second ExtInfo later to authenticated clients to avoid disclosing extension support to unauthenticated clients. Each ExtInfo message contains several extension entries. Negotiation requirements are defined on a per-extension level.
RFC 8308 defines an initial set of four protocol extensions, and vendors have proposed and implemented additional extensions. We detail those relevant to our attacks here and describe the others in Appendix B.
@@
publickey-hostbound@openssh.com [32, 31] is a server-side extension to advertise support for host-bound public key authentication, which deviates from public key authentication by also covering the server’s host key. This allows the enforcement of per-key restrictions when allowing remote servers to access local secret keys (i.e., when using SSH Agent).
ping@openssh.com [32] is a server-side extension to advertise support for a transport level ping message similar to the Heartbeat extension in TLS [41].
@@ 5.2 Extension Downgrade Attack
We now show how the prefix truncation attack can be applied to delete the ExtInfo message sent by the client, server, or both parties without either noticing. Our attack differs depending on the encryption mode. For ChaCha20-Poly1305, we can use the basic attack strategy. For CBC-EtM, we show two strategies to generate additional messages in the secure channel, so that the attacker can use the “evasively corrupt” outcome of Scenario 2 in Section 4.3.
@@
(b) Extension Downgrade Attack for CBC-EtM: The MitM injects Unknown before the NewKeys is sent by the client. As the server already sent NewKeys, the provoked Unimplemented message will be sent within the secure channel after ExtInfo. The corrupted Unimplemented message has a significant probability of being ignored (see Scenario 2 in Section 4.3).
@@
The downgrade attack for ChaCha20-Poly1305 against the client is depicted in 3(a). It is identical to the single message prefix truncation attack from Section 4.2, with ExtInfo now taking the place of SC1 in Figure 2. If the attack should be directed against the server instead, a $(0,1)$-prefix truncation attack should be performed instead. This allows an attacker to delete any ExtInfo sent immediately after NewKeys.
While the server may send a second ExtInfo just before signaling successful client authentication, stripping the ExtInfo message sent after NewKeys renders most publicly specified extensions unusable. This is because they are either scoped to the authentication protocol, sent by the client only, or must be sent by both parties to take effect. Solely the ping@openssh.com extension may be sent in the second ExtInfo to enable keystroke timing countermeasures inside the connection protocol. However, OpenSSH 9.5 does not implement any facility to send a second extension negotiation message. As we show in Section 7, extensions scoped to the authentication protocol are the most common among SSH servers on the internet by a significant margin.
@@
In 3(b), we show how the attack can also work with CBC-EtM. Suppose an attacker injects an Unknown message to the server after the server sends NewKeys and ExtInfo, but before the client’s NewKeys message (and also injects Unknown to the client to realign sequence numbers). In that case, the server sends the response Unimplemented as the second message in the secure channel immediately after the ExtInfo message. The attacker now wants to remove two messages from the channel and can benefit from the “evasively corrupt” in Scenario 2 in Section 4.3. The attacker removes ExtInfo from the secure channel, which causes the decryption of the first block of Unimplemented to be pseudo-random. Because Unimplemented messages are relatively small ($\ell=16$ for AES), the upper estimate for the success probability is only $11\cdot 213\cdot 2^{-16}\approx 0.0358$.
@@
We successfully evaluated the attack on ChaCha20-Poly1305 and CBC-EtM against OpenSSH 9.5p1 and PuTTY 0.79 clients, connecting to OpenSSH 9.4p1 (Unknown only) and 9.5p1. For CBC-EtM, our success rate in practice was 0.0003 (OpenSSH) resp. 0.03 (PuTTY), improved to 0.0074 (OpenSSH) resp. 0.8383 (PuTTY) when sending Ping instead of Unknown.
@@ 6 Message Injection Attacks on AsyncSSH
Going beyond the SSH specifications, we now demonstrate how prefix truncation attacks can also be used to exploit implementation flaws. Specifically, we target AsyncSSH, [^2] https://github.com/ronf/asyncssh an SSH implementation for Python with an estimated 60k daily downloads. [^3] https://pypistats.org/packages/asyncssh We present two attacks that exploit weaknesses in the handling of unauthenticated messages during the handshake. These attacks are enabled by prefix truncation and sequence number manipulation.
Note that we describe these attacks only for ChaCha20-Poly1305. Adjusting them for CBC-EtM is straightforward, injecting appropriate Ignore and Unknown messages, but requires some of the advanced techniques described in Appendix A. These advanced techniques only work against some SSH implementations.

### v2

@@ 4.4.2 Affected And Perfectly Exploitable
ChaCha20-Poly1305 [34] directly uses the sequence number in its internal key stream derivation, which makes it vulnerable to our prefix truncation attack. All messages following the truncated prefix are decrypted to their original plaintext because the integrity check of the AEAD cipher is done over the ciphertext and the sequence number, which the attacker has manipulated to match. Under the assumption that the attacker can correctly guess the packet length, the prefix truncation attack always succeeds.
Note that the fault is not with ChaCha20-Poly1305 as an AEAD encryption scheme but with its integration into the SSH secure channel construction.
4.4.3 Affected With Limited Exploitability
@@
With CTR-EtM, the MAC is computed over the unencrypted length, the sequence number, and the ciphertext. So, removing some packets from the beginning of the channel does not cause a MAC failure, and cryptographically, the attack succeeds. However, CTR uses a block counter initialized to $\mathrm{IV}_{\mathrm{KDF}}$, which increments after each block. After prefix truncation, the key stream is desynchronized, so all following ciphertexts are decrypted as pseudorandom packets. Each corrupted packet has a significant probability of causing a critical failure, eventually stopping our attack.
@@
For CTR-EtM, prefix truncation of $k$ blocks (which exactly contain one or more messages) provides a very limited decryption oracle on the ciphertext $c_{1},\ldots,c_{k}$ where $c_{i}:=\mathrm{Enc(\mathrm{IV}_{\mathrm{KDF}}+i)}\oplus p_{i},1\leq i\leq k$. After deleting the first $k$ blocks, MAC verification for the following message of length $l$ blocks will succeed because the length, sequence number, and ciphertext are correct. The blocks $c_{k+1},\ldots,c_{k+l}$ will be decrypted as $p^{\prime}_{j}:=\mathrm{Enc(\mathrm{IV}_{\mathrm{KDF}}+j)}\oplus c_{k+j},1\leq j\leq l$, and processed as a pseudorandom SSH message SC1’. Due to format oracle side channels in SSH at the BPP layer, e.g., the padding length, but also at the protocol layer, e.g., if a message is ignored or triggers a response, the attacker can get some information about the bits in $p^{\prime}_{j}$. This reveals information about the first $l$ key stream blocks, and thus also about $p_{1},\ldots,p_{l}$, potentially leaking confidential information like passwords in user authentication. If processing SC1’ does not cause a critical failure, the attack can even continue, revealing more about the following key stream and, thus, plaintext. Exploiting this requires a careful study of format oracles in SSH, which is outside the scope of this work.
@@
Recall that the decryption of the first block is $p_{1}:=\mathrm{Dec}(c_{1})\oplus\mathrm{IV}_{\mathrm{KDF}}$, and for block $i$, it is $p_{i}:=\mathrm{Dec}(c_{i})\oplus c_{i-1}$. We assume the attacker uses prefix truncation to remove blocks $c_{1},\ldots,c_{k}$. The following block $c_{k+1}$ will now be decrypted as $p^{\prime}_{1}:=\mathrm{Dec}(c_{k+1})\oplus\mathrm{IV}_{\mathrm{KDF}}$. We are interested in how SSH implementations process the resulting pseudorandom block $p^{\prime}_{1}$ as the first block in the decrypted packet. Intuitively, it should result in a corrupted packet that causes a critical failure.222 Similarly to CTR-EtM, any format oracle side channel for $p^{\prime}_{1}$ reveals a relationship between $\mathrm{IV}_{\mathrm{KDF}}$ and $p_{k+1}$ via $\mathrm{IV}_{\mathrm{KDF}}\oplus p_{k+1}=c_{k}\oplus p^{\prime}_{1}$, which is a marginal information leak for the (secret) IV given information on $p_{k+1}$, and vice versa. Again, we do not explore this further here.
Surprisingly, there is a significant probability that the attack can continue, although it is highly implementation-dependent. For a corrupted packet, there are four possible outcomes:
@@
    Critically Corrupt: If corruption is detected at the BPP or application level, e.g., if a length field exceeds the packet length, the connection should be closed.
@@
    Evasively Corrupt: If the packet is well-formed (i.e., has valid padding length) but has an unrecognized message ID, an Unimplemented response must be sent, and the connection continues normally [33, Sec. 11.4].
  4. 4.
    Any other case not covered above, in particular, recognized messages different from the original.
Clearly, the first outcome stops any attack from going forward. However, the second, third, and fourth outcomes may be beneficial for the attacker. We will now present two instructive scenarios for outcomes two and three, and estimate the success probability of an attack relying on that outcome. Later, we will verify these estimates experimentally.
@@
Although ServiceAccept may be a lucky case (for the attacker), there are structural reasons for this result: First, SSH messages are often short and can be smaller than a single block. Second, the padding is random and cannot be verified. Third, some messages have redundant fields that implementations ignore (e.g., the service name above).
We experimentally verified that OpenSSH, Dropbear, PuTTY, and libssh allow empty ServiceAccept messages from the server, enabling this attack. At the same time, AsyncSSH is strict by requiring the correct service name.
@@
In this scenario, we assume the attacker wants to remove the first $N>1$ messages and preserve all the following messages perfectly. Then, the attacker can use prefix truncation to delete the first $N-1$ messages and take a bet on the $N$-th message to be evasively corrupt.
@@
Because the message is well-formed, it is not rejected at the BPP layer. Furthermore, because the message is unrecognized, the peer must respond with Unimplemented and otherwise ignore it [33, Sec. 11.4], so our attack succeeds.
@@
The probability that a packet is unrecognized depends on the size of the set $U$ of unrecognized message IDs in the implementation. The attack requires at least one unknown message ID. Through source code review, we identified 43 IDs that are in active use, so we estimate up to 213 unknown message IDs out of $2^{8}$.
In total, we estimate a success probability of $\mathrm{min}(252,\ell-5)\cdot|U|\cdot 2^{-16}$. Assuming a block size of at least 128-bit (i.e., $\ell\geq 16$), we estimate that the success probability of this attack is between $11\cdot 2^{-16}\approx 0.0002$ ($\ell_{\mathrm{min}}=16,|U_{\mathrm{min}}|=1$) and $252\cdot 213\cdot 2^{-16}\approx 0.8190$ ($\ell_{\mathrm{max}}\geq 252,|U_{\mathrm{max}}|=213$) for vulnerable implementations. Our experiments show success probabilities from 0.0003–0.8383, in good agreement with our analysis (Section 5.2). Increasing the block size increases the lower bound, while the upper bound stays the same.
@@ 5 Breaking SSH Extension Negotiation
While the fact that BPP does not implement a secure channel is troublesome enough, exploiting this vulnerability requires an analysis of the SSH protocol after the handshake, i.e., the SSH authentication protocol.
As our attack achieves prefix truncation, it is natural to ask which SSH messages can occur at the beginning of a secure channel. Historically, the first messages exchanged are ServiceRequest and ServiceAccept. Removing either causes the connection to go stale, as the client will not begin the user authentication. Then, our attack, while cryptographically successful, fails at the application layer.
However, the SSH Extension Negotiation mechanism [9] introduces a new message, ExtInfo, which can occur immediately after NewKeys as the first message on the secure channel. Some of the extensions that can be negotiated are security-relevant, providing an attack surface for our prefix truncation attack and raising its impact.
@@ 5.1 SSH Extension Negotiation
Support for extension negotiation is signaled as part of the KexInit message. The structure of the message is not altered, and the reserved field is not used to avoid compatibility issues. Instead, each peer may include an indicator name within the list of key exchange algorithms. The indicator name differs depending on the role of the peer (ext-info-c vs. ext-info-s) to avoid accidental negotiation.
Whenever a peer signals support for extension negotiation, the other side may send an ExtInfo message as the first message after NewKeys. Additionally, the server can send a second ExtInfo later to authenticated clients to avoid disclosing extension support to unauthenticated clients. Each ExtInfo message can contain several extension entries. Negotiation requirements are defined on a per-extension level.
RFC 8308 defines an initial set of four protocol extensions, and vendors have proposed and implemented additional extensions. We detail those relevant to our attacks here. [^3] We excluded the following extensions because we consider them unrelated to our attacks: no-flow-control, delay-compression, elevation, global-requests-ok, ext-auth-info
@@
publickey-hostbound@openssh.com [36, 35] is a server-side extension to advertise support for host-bound public key authentication, which deviates from public key authentication by also covering the server’s host key. This allows the enforcement of per-key restrictions when generating the signature outside the SSH client (i.e., when using SSH Agent).
ping@openssh.com [36] is a server-side extension to advertise support for a transport-level ping message similar to the Heartbeat extension in TLS [46].
@@ 5.2 Extension Downgrade Attack
We now show how the prefix truncation attack can be applied to delete the ExtInfo message sent by the client, server, or both parties without either noticing. Our attack differs depending on the encryption mode. For ChaCha20-Poly1305, we can use the basic attack strategy. For CBC-EtM, we show two strategies to generate additional messages in the secure channel so that the attacker can use the “evasively corrupt” outcome of Scenario 2 in Section 4.4.3.
Impact
@@
(b) Extension Downgrade Attack for CBC-EtM: The MitM injects Unknown before the NewKeys is sent by the client. As the server already sent NewKeys, the provoked Unimplemented message will be sent within the secure channel after ExtInfo. The corrupted Unimplemented message has a significant probability of being ignored (see Scenario 2 in Section 4.4.3).
@@
The downgrade attack for ChaCha20-Poly1305 against the client is depicted in 5(a). It is identical to the single message prefix truncation attack from Section 4.2, with ExtInfo now taking the place of SC1 in Figure 2. If the attack should be directed against the server instead, a $(0,1)$-prefix truncation attack should be performed. This allows an attacker to delete any ExtInfo sent immediately after NewKeys.
While the server may send a second ExtInfo just before signaling successful client authentication, stripping the ExtInfo message sent after NewKeys renders most publicly specified extensions unusable. This is because they are either scoped to the authentication protocol, sent by the client only, or must be sent by both parties to take effect. Solely the ping@openssh.com extension may be sent in the second ExtInfo to enable keystroke timing countermeasures inside the connection protocol. However, OpenSSH 9.5 does not implement any facility to send a second extension negotiation message. As shown in Section 7, extensions scoped to the authentication protocol are the most common among SSH servers on the internet by a significant margin.
@@
In 5(b), we show how the attack can also work with CBC-EtM. Suppose an attacker injects an Unknown message to the server after the server sends NewKeys and ExtInfo but before the client’s NewKeys message (and also injects Unknown to the client to realign sequence numbers). In that case, the server sends the response Unimplemented as the second message in the secure channel immediately after the ExtInfo message. The attacker now wants to remove two messages from the channel and can benefit from the “evasively corrupt” in Scenario 2 in Section 4.4.3. The attacker removes ExtInfo from the secure channel, which causes the decryption of the first block of Unimplemented to become pseudorandom. Because Unimplemented messages are relatively small ($\ell=16$ for AES), the upper estimate for the success probability is only $11\cdot 213\cdot 2^{-16}\approx 0.0358$.
@@
We successfully evaluated the attack in 10,000 trials on ChaCha20-Poly1305 and CBC-EtM against OpenSSH 9.5p1 and PuTTY 0.79 clients, connecting to OpenSSH 9.4p1 (Unknown only) and 9.5p1. For CBC-EtM, our success rate in practice was 0.0003 (OpenSSH) resp. 0.0300 (PuTTY), improved to 0.0074 (OpenSSH) resp. 0.8383 (PuTTY) when sending Ping instead of Unknown.
@@ 6 Message Injection Attacks on AsyncSSH
Going beyond the SSH specifications, we now demonstrate how prefix truncation attacks can also be used to exploit implementation flaws. Specifically, we target AsyncSSH, [^4] https://github.com/ronf/asyncssh an SSH implementation for Python with an estimated 60k daily downloads. [^5] https://pypistats.org/packages/asyncssh We present two attacks that exploit weaknesses in handling unauthenticated messages during the handshake. These attacks are enabled by prefix truncation and sequence number manipulation.
Note that we describe these attacks only for ChaCha20-Poly1305. Adjusting them for CBC-EtM is straightforward, injecting appropriate Ignore and Unknown messages, but requires some of the advanced techniques described in Section 4.1. These advanced techniques only work against some SSH implementations.